OpenStudy (anonymous):

Determine whether the sequence converges or diverges. If it converges, give the limit. 60, -10, 5/3, -5/18, ...

OpenStudy (zarkon):

what do you think?

OpenStudy (anonymous):

it converges i guess maam

OpenStudy (zarkon):

looks like \(\Large a_n=\frac{60}{(-6)^{n-1}}\) for \(n\in \mathbb{N}\)

OpenStudy (anonymous):

So my choices are: Diverges Converges; 11100 Converges; 72 Converges; 0

OpenStudy (zarkon):

compute \[\Large \lim_{n\to\infty}\frac{60}{(-6)^{n-1}}\]

OpenStudy (anonymous):

I don't know what any of that means.. :(

OpenStudy (zarkon):

what class is this for?

OpenStudy (anonymous):

converges at 0 i guess @Zarkon am i correct?

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

converges to 0

OpenStudy (anonymous):

It's precalculus

OpenStudy (zarkon):

ic

OpenStudy (anonymous):

Thanks guys!!